### v2

# Eliciting In-context Retrieval and Reasoning for Long-context Large Language Models

Abstract
Recent advancements in long-context language models (LCLMs) promise to transform Retrieval-Augmented Generation (RAG) by simplifying pipelines. With their expanded context windows, LCLMs can process entire knowledge bases and perform retrieval and reasoning directly – a capability we define as In-Context Retrieval and Reasoning (ICR2). However, existing benchmarks like LOFT often overestimate LCLM performance by providing overly simplified contexts. To address this, we introduce ICR2, a benchmark that evaluates LCLMs in more realistic scenarios by including confounding passages retrieved with strong retrievers. We then propose three methods to enhance LCLM performance: (1) retrieve-then-generate fine-tuning, (2) retrieval-attention-probing, which uses attention heads to filter and de-noise long contexts during decoding, and (3) joint retrieval head training alongside the generation head. Our evaluation of five well-known LCLMs on LOFT and ICR2 demonstrates significant gains with our best approach applied to Mistral-7B: +17 and +15 points by Exact Match on LOFT, and +13 and +2 points on ICR2, compared to vanilla RAG and supervised fine-tuning, respectively. It even outperforms GPT-4-Turbo on most tasks despite being a much smaller model.
Eliciting In-context Retrieval and Reasoning for
Long-context Large Language Models
Yifu Qiu1thanks:   Work done while the author was an intern at Apple., Varun Embar2, Yizhe Zhang2, Navdeep Jaitly2, Shay B. Cohen1, Benjamin Han2 1 University of Edinburgh, 2 Apple {yifu.qiu, scohen}@ed.ac.uk , {v_embar, yizhe_zhang, njaitly, ben.b.han}@apple.com
1 Introduction
The ability of large language models to process long contexts has significantly expanded their applicability across various domains, including book-level information retrieval (Ding et al., 2024; Jin et al., 2024), summarization (Kim et al., 2024; Saxena and Keller, 2024; Qiu et al., 2023), and question answering (Liu et al., 2024; Wang et al., 2024a). They also enable more complex tasks, such as agent trajectory modeling and planning (Zhao et al., 2024; Zhang et al., 2024b), video captioning (Xue et al., 2024; Zhang et al., 2023), and text-to-video generation (Wang et al., 2024b; Lin et al., 2023). Recent advancements in long-context language models (LCLMs) hold particular promise for reshaping the Retrieval-Augmented Generation (RAG) paradigm (Lee et al., 2024; Li et al., 2024). With their expanded context windows, LCLMs reduce reliance on complex pipelines required by previous context-length limitations and simplify knowledge updates by allowing context modifications. For instance, LCLMs can accommodate entire knowledge bases within their context windows, effectively serving as working memory to be used for new queries.
Achieving this goal requires LCLMs to effectively retrieve and reason within their “contextual knowledge base,” a capability we define as In-Context Retrieval and Reasoning (ICR2). However, existing benchmarks often fail to accurately evaluate this capability. For example, Needle-in-a-Haystack (NIAH; Kamradt 2023) is a popular test to determine whether a model can retrieve a “needle” (a specific fact or statement) randomly inserted into a “haystack” (a corpus). Yet, the semantic discontinuity between the needle and the haystack can unintentionally reveal the needle’s location, making the task overly simple. LOFT (Lee et al., 2024), the first large-scale benchmark for evaluating retrieval and reasoning within contextual knowledge bases, uses human-annotated relevant documents as the needle, and fills the haystack with randomly sampled documents from an external knowledge base. However, this random sampling results in a context with virtually no confounding information that is relevant but misleading, causing LOFT to significantly overestimate LCLM performance.
To bridge this evaluation gap, we introduce ICR2, a novel and challenging benchmark designed to assess LCLMs under more realistic conditions. ICR2 builds upon KILT (Petroni et al., 2021), a comprehensive knowledge base sourced from Wikipedia. Unlike LOFT, which relies on random sampling, ICR2 uses strong retrievers to select challenging confounding documents, creating a more difficult “haystack.” Experimental results reveal that current LCLMs struggle on ICR2, with exact match rates dropping by up to 51% compared to evaluations on LOFT. These findings underscore the significant challenges LCLMs face in accurately performing in-context retrieval in realistic scenarios.
We next explore improving LCLMs’ in-context retrieval and reasoning capabilities. While RAG demonstrates strong results, it remains hindered by complex multi-stage pipelines. Encouraged by recent studies such as LLM2Vec (BehnamGhader et al., 2024; Ma et al., 2024) showing that LCLMs can be effectively adapted for accurate retrieval tasks, offering a more natural approach by enabling joint optimization of both retrieval and generation steps, we propose three approaches: (1) Retrieve-then-generate fine-tuning: Inspired by the distillation of step-by-step reasoning abilities (Shridhar et al., 2023; Hsieh et al., 2023), we train LCLMs to frame tasks as two-hop reasoning chains. In this formulation, LCLMs first retrieve relevant information from the context and then generate the final responses. (2) Retrieval-attention probing: At inference time, we probe attention heads activated for in-context retrieval (Wu et al., 2024) and use their top predictions to filter out confounders from lengthy contexts. (3) Joint retrieval head training: We introduce an architectural modification to equip LCLMs with a dedicated retrieval head, enabling joint optimization of retrieval and generation.
We conduct extensive experiments using five LCLMs on both LOFT and our ICR2 benchmark. We compare our methods against baselines including Vanilla RAG, Closed-book, Oracle RAG, and supervised fine-tuning (SFT). Notably, our best approach, applied to Mistral-7B with a 32K token limit, has the best performance across the tasks. It outperforms Vanilla RAG and SFT baselines by an average of +17 and +15 points measured by Exact Match on LOFT, and by +13 and +2 points on ICR2, respectively. The approach achieves performance comparable to the state-of-the-art GPT-4, despite using only 7B parameters. We also provide in-depth analyses of our approaches.
Our key contributions are as follows:
  • We introduce ICR2, a realistic and challenging benchmark for evaluating the in-context retrieval and reasoning capabilities of long-context language models (LCLMs), demonstrating that existing benchmarks overestimate model performance.
  • We explore three novel methods to enhance LCLMs ranging from supervised fine-tuning, inference-time approach, to model architecture modification.
  • Our best approach, applied to a small LCLM, tightens the performance gap with the Oracle RAG while beating the other baselines. It even matches GPT-4 on both in-domain (ICR2) and out-of-domain benchmarks (LOFT), albeit with a much smaller model size.
2 Related Work
Long-context large language models (LCLMs) have garnered significant attention for their ability to process extended sequences. Models like Longformer (Beltagy et al., 2020) and BigBird (Zaheer et al., 2020) introduced sparse attention mechanisms to efficiently handle long context. Scaling efforts, exemplified by GPT-4 (Achiam et al., 2023), underscore the importance of expanding context windows for the long-context tasks. Data engineering methods (Fu et al., 2024; Xiong et al., 2024; Jin et al., 2024), expanding positional encoding (Ding et al., 2024), and the parameter-efficient fine-tuning (Chen et al., 2024) have also been proven to be effective.
LCLMs hold the promise in reshaping RAG (Achiam et al. 2023; Jiang et al. 2023; Yang et al. 2024; Abdin et al. 2024). By replacing static knowledge bases with the contextual one, this new paradigm simplifies the deployment by discarding intermediate components while enabling the direct updates on LCLM’s knowledge (Lee et al., 2024).
Unlike the existing works, we systematically evaluate the use of LCLMs for scenarios where knowledge bases are directly placed as the context with a novel benchmark, ICR2. ICR2 is different with the popular long-context benchmark such as (Bai et al., 2024) by emphasizing in the contextual confounders. We also show that with targeted enhancements, small-scale LCLMs can achieve performance comparable to state-of-the-art models.
3 Are LCLMs Competent for RAG?
3.1 LLMs Are Sensitive to the Confounders
Figure 1: Task-specific performance of five LCLMs on 32K ICR2 test sets with varying confounding ratios. For all tasks, a higher value of task-specific metric indicates a better performance.
LOFT (Lee et al., 2024) introduces a Corpus-in-Context (CiC) approach for retrieval-augmented generation, integrating a large-scale external knowledge base directly into the LLM’s context. However, LOFT operates under the assumption that the context is free of confounders. In practice, a corpus often contains confounders – documents related to the query but potentially leading to incorrect answers. For example, given the query “Who is the 44th U.S. President?”, the corpus might include documents about the other presidents, such as “The 45th U.S. President is Donald Trump”, which could mislead the LLM. Consequently, LOFT’s design reduces the complexities of using real-world corpora for RAG, potentially overestimating the performance of LCLMs.
We show that LCLMs are indeed sensitive to the confounders missed in LOFT. We construct multiple test sets with varying confounding ratios$\{0,25\%,50\%,75\%,100\%\}$ – and evaluate the LCLMs under zero-shot settings. The confounding ratio $p$ denotes the proportion of confounding context that are selected by retrievers while filtering out the gold provenance, with the remaining, i.e., $(100\%-p)$, randomly sampled from an external knowledge base. At $p=0$, all confounding passages in the contextual knowledge base are randomly sampled, equivalent to the setup in LOFT (Lee et al., 2024). Conversely, at $p=1$, all confounders are selected by the retrievers.
We evaluate five LCLMs with a context length of at least 32K tokens: Phi-3-7B (Abdin et al., 2024), Qwen-2-1.5B/7B (Yang et al., 2024), Mistral-003-7B (Jiang et al., 2023), LLaMA-3-Instruct-8B (Dubey et al., 2024) and GPT-4-Turbo (Achiam et al., 2023). Our findings in Figure 1, demonstrates that LCLM performance is highly sensitive to confounders, with performance generally degrading as the confounding ratio increases. This indicates that confounders in $P^{-}$ obtained via retrievers, as in ICR2, pose greater challenges for the models compared to those randomly sampled in LOFT. These retriever-selected passages are relevant to the queries but fail to lead to the correct answers, making them particularly difficult for LCLMs to handle. Consequently, benchmarks like LOFT that overlook these “strong” confounders risk overestimating model performance, as confounders are prevalent in real-world scenarios.
3.2 ICR2 Benchmark
Figure 2: Construction pipeline for ICR2. Based on KILT, we use five strong retrievers to retrieve confounders to serve as the “haystack”, while the human-annotated contexts are used as the “needle”.
To address this limitation, we propose an alternative benchmark, ICR2, which leverages strong retrievers to identify and incorporate these confounders into the contextual corpus used in CiC, providing a more realistic and challenging evaluation framework. We choose KILT (Petroni et al., 2021), a comprehensive suite of benchmarks designed for knowledge-intensive NLP tasks, to be our external knowledge base (Figure 2). KILT covers tasks such as question answering (Kwiatkowski et al., 2019; Yang et al., 2018), fact verification (Thorne et al., 2018), and dialogue completion (Dinan et al., 2018), all paired with a single Wikipedia snapshot. Each KILT instance $\langle q,a,P^{+}\rangle$ consists of a query $q$, the reference answer $a$, and its provenances $P^{+}=\{p_{1},\dots,p_{m}\}$, which is a set of relevant Wikipedia pages and specific locations within them that support the answer.
[…]
Query: Where are the giant redwoods located in California? Reference Answer: Humboldt County Contextual Knowledge Base: ID: PSG001 Title: 2010–11 California Golden Bears Men’s Basketball Team Context: The 2010–11 California Golden Bears men’s basketball team represented the University of California, Berkeley in the 2010–11 NCAA Division I men’s basketball season. ID: PSG002 Title: Redwood National and State Parks Context: Scenes set on the forest moon Endor in Star Wars were filmed in the Tall Trees Redwood Grove in the northern part of Humboldt County, though the majority of filming was in private and public forests near the town of Smith River … … [90 context are omitted for the simplicity] … ID: PSG100 Title: Malibu Grand Prix Context: … and San Antonio, Texas. Palace operates additional locations in Los Angeles, California and Dallas, Texas. The Redwood City, CA location closed on August 18, 2013 and the San Antonio location closed on September 7, 2015.
Table 1: An example ICR2 instance. We show the query with the reference answer, and the constructed contextual knowledge base. We highlight the true provenance in blue and the confounding passages in red.
We show the pipeline for ICR2 creation in Figure 2. To eliminate bias in the retrievers’ preferences during the construction process, we pool the Top-$K$ predictions from all retrievers and uniformly sample passages from top to bottom until reaching the maximum allowable size for the contextual knowledge base. The retrievers used include Contriever (Izacard et al., ), DPR (Karpukhin et al., 2020), DrQA (Chen et al., 2017), BM25 (Robertson and Jones, 1976; Robertson et al., 2009), and BLINK (Wu et al., 2020). For passage-level retrievers such as Contriever, BM25, and DPR, as input we chunk all documents in KILT’s knowledge base and exclude passages overlapping with any annotated provenance in $P^{+}$. We also set $K=200$ for the 32K version of ICR2. For document-level retrievers like DrQA and BLINK, we set $K=20$, and chunk retrieved documents using the same process as above. For all retrievers, we exclude retrieved passages overlapping with $P^{+}$ or containing exact substring matches with $a$, and the remaining passages are collected to form the confounding provenance set, $P^{-}$. The contextual corpus $C$ is then constructed by concatenating $P^{+}$ with $P^{-}$ up to the maximum length, i.e., $C=[P^{+};P^{-}]$. Finally, we shuffle all items in $C$ to remove position bias.
Table 2 compares statistics between LOFT and our benchmark ICR2. An example of an ICR2 instance is provided in Table 1, and the rag_prompt template used is detailed in Table 8 in Appendix A.
Task #CTX #Tokens #Prov
LOFT NQ 215 28,911 1
HPQA 276 28,924 1.98
MUSIQ. 207 28,806 1.67
ICR2 NQ 202 20,441 1
HPQA 202 20,818 2.09
FEVER 202 21,136 1
WoW 202 21,068 1
Table 2: Statistics for LOFT and our dataset ICR2: #CTX is to the average number of passages per query in the contextual knowledge base, #Tokens is the average length of the CiC prompt, and #Prov. is the average number of provenance (positive passages) per query.
3.3 State of LCLMs Performance on ICR2
LOFT ICR2 Avg-All
NQ HotpotQA MUSIQUE Avg NQ HotpotQA FEVER WoW Avg
Closed-book
Qwen-2-1.5B-32K 0.21 0.22 0.03 0.15 0.29 0.24 0.62 0.15 0.32 0.25
Qwen-2-7B-32K 0.31 0.30 0.07 0.23 0.35 0.30 0.71 0.15 0.38 0.31
Mistral-2-7B-32K 0.46 0.38 0.05 0.30 0.49 0.38 0.73 0.14 0.43 0.38
Phi-3-7B-128K 0.39 0.35 0.09 0.28 0.46 0.32 0.73 0.13 0.41 0.35
LLaMA-3-instruct-8B 0.41 0.19 0.01 0.20 0.44 0.19 0.53 0.14 0.32 0.27
GPT-4-turbo 0.58 0.57 0.28 0.48 0.61 0.50 0.85 0.16 0.53 0.51
Vanilla RAG
Qwen-2-1.5B-32K 0.61 0.43 0.12 0.39 0.35 0.32 0.78 0.12 0.39 0.39
Qwen-2-7B-32K 0.79 0.61 0.29 0.56 0.57 0.47 0.88 0.13 0.51 0.53
Mistral-2-7B-32K 0.64 0.62 0.27 0.51 0.53 0.50 0.81 0.13 0.49 0.50
Phi-3-7B-128K 0.76 0.68 0.41 0.62 0.60 0.50 0.90 0.14 0.53 0.57
LLaMA-3-instruct-8B 0.45 0.56 0.16 0.39 0.53 0.45 0.83 0.15 0.49 0.45
GPT-4-turbo 0.85 0.78 0.51 0.71 0.67 0.64 0.92 0.15 0.59 0.65
Oracle RAG
Qwen-2-1.5B-32K 0.80 0.76 0.44 0.67 0.77 0.65 0.86 0.15 0.61 0.63
Qwen-2-7B-32K 0.88 0.81 0.56 0.75 0.83 0.80 0.91 0.17 0.68 0.71
Mistral-2-7B-32K 0.89 0.81 0.41 0.70 0.83 0.81 0.94 0.18 0.69 0.70
Phi-3-7B-128K 0.90 0.85 0.63 0.79 0.87 0.81 0.91 0.14 0.68 0.73
LLaMA-3-instruct-8B 0.81 0.77 0.60 0.73 0.80 0.74 0.86 0.19 0.65 0.68
GPT-4-turbo 0.88 0.87 0.72 0.82 0.79 0.82 0.94 0.18 0.68 0.74
Table 3: Performance evaluation of six LCLMs on LOFT and ICR2 benchmarks. All models benefit from the Vanilla RAG approach, but a gap remains between vanilla and oracle RAG performance.
We evaluate five LCLMs that support 32K or longer context length using LOFT (Lee et al., 2024) and ICR2 benchmarks, with each one tested with input up to 32K tokens. We include three baselines for comparison: (1) Vanilla RAG, where the CiC prompt from Lee et al. (2024) is used; (2) Closed-book, where the entire context ($C$) is removed to evaluate LCLM performance based solely on parametric knowledge; and (3) Oracle RAG, which includes only the ground-truth relevant contexts $Z^{*}$ in $C$ to estimate an upper-bound performance.
Table 3 presents our results. Overall, we observe that the Vanilla RAG significantly outperforms the closed-book baseline for most models, except for Qwen-2-1.5B on the WoW dataset. This trend persists even with the state-of-the-art GPT-4-Turbo, suggesting that external knowledge remains crucial for knowledge-intensive tasks despite the large parametric capacity of LCLMs. Furthermore, RAG performance on ICR2 lags notably behind that on LOFT, reflecting ICR2’s greater complexity due to its more realistic context construction. Lastly, the large performance gap between RAG and Oracle setups on both benchmarks underscores the impact of confounding information, which hampers accurate retrieval and response generation in LCLMs.
4 Eliciting ICR2 for LLMs
4.1 Retrieve-then-generate Fine-tuning
Compared to the standard supervised fine-tuning where a model is asked to generate a Direct Answer (DA) to a given query, our first proposal, retrieve-then-generate, is a two-step process: the model first retrieves relevant passages from context and then generates the final answer based on the context and retrieved passages, all in one decoding pass. Formally, we train an LCLM to optimize the objective,
$p(y\mid q,c)=\sum_{z_{i}\in Z}p(y\mid q,c,z_{i})p(z_{i}\mid q,c),$ (1)
where $q,y$ is the query and target, respectively, $c$ is the contextual knowledge base, and $Z$ is the collection of all relevant passages necessary for answering $q$. This objective can be easily integrated into the next-token prediction task trained with the maximum likelihood estimation, just by sequentially executing the retrieval and generation in a single forward pass. The overall loss is given as
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}\left(\log p(Z_{i}^{*}|q_{i},c_{i})+\log p(y_{i}^{*}|Z_{i}^{*},q_{i},c_{i})\right),$ (2)
where $N$ is the number of training samples, and $Z^{*}_{i}=\{z^{*}_{i,1},z^{*}_{i,2},\dots,z^{*}_{i,|Z^{*}_{i}|}\}$ is the collection of all relevant passages for the query $q_{i}$ and contextual knowledge base $c_{i}$.
We implement two variants of retrieve-then-generate fine-tuning, both using the special tokens <RETRIEVAL> and </RETRIEVAL> to delineate the retrieval step. In the first, Retrieve-Then-Answer (RTA), the model copies relevant passages verbatim during retrieval. In the second, Cite-Context-ID (CCI), the model generates only the IDs of the relevant passages. Figure 3 illustrates the templates for these variations.
Figure 3: Templates for Direct Answer (DA), Retrieve-Then-Answer (RTA) and Cite-Context-ID (CCI).
4.2 Retrieval Attention Probing
Our second proposal, Retrieval Attention Probing (RAP), is an inference-time approach compatible with LCLMs without requiring re-training. Building on Wu et al.’s findings that specific attention heads are highly active during retrieval tasks (e.g., NIAH), RAP utilizes these retrieval-focused attention heads for context filtering before generating responses. For each attention head $h$ and query $q$, we track the Top-$M$ attention scores $A_{h}^{M}(q)$. $C_{h}^{M}(q)$, representing the $M$ passages corresponding to $A_{h}^{M}(q)$, are then selected, and the hit rate for head $h$ is calculated as follows:
$\text{HitRate}_{h}$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{|C_{h}^{M}(q)\cap Z^{*}(q)|}{|Z^{*}(q)|},$ (3)
where $N$ is the number of validation samples, and $Z^{*}(q)$ is the set of all relevant contexts given query $q$. Finally we select $Q$ heads with the top hit rates to be the retrieval heads as follows,
$\mathcal{H}_{\text{ret}}=\underset{\mathcal{H}\subseteq\{1,\dots,H\},|\mathcal{H}_{\text{ret}}|=Q}{\arg\max}\sum_{h\in\mathcal{H}}\text{HitRate}_{h},$ (4)
where $H$ is the number of attention heads, and $\mathcal{H}_{\text{ret}}$ is the set of $Q$ attention heads with the top hit rates selected as the retrieval heads for context filtering.
Note that our definition of retrieval heads differs from (Wu et al., 2024) in two key ways: (1) Instead of focusing solely on the Top-$1$ passages, we allow each head to retain the Top-$M$, which is particularly important for multi-hop reasoning tasks requiring multiple passages for reasoning. (2) We evaluate attention heads using the retrieval hit rate, a more direct metric for our downstream tasks.
During inference, we union all passages selected by all retrieval heads, $\mathcal{H}_{\text{ret}}$,
$C^{*}=\bigcup_{h\in\mathcal{H}_{\text{ret}}}\underset{c\in C}{\text{Top-}M}\,\alpha_{h}(c),$ (5)
where $\alpha_{h}(c)$ is the attention score of head $h$ on passage $c$, and $C^{*}$ is the set of the Top-$M$ selected passages from all heads $\mathcal{H}_{\text{ret}}$. We use $C^{*}$ to form a new CiC-style prompt and proceed to generating the final response. Since $|C^{*}|\ll|C|$, the final decoding is actually performed on a filtered contextual knowledge base with a much smaller length.
4.3 Joint Retrieval Head Training
Our final proposal introduces a dedicated retrieval head to the LCLM model architecture. During inference, the model first uses the head to identify relevant passages, after which the generation head decodes a response conditioned on the retrieved content. During training, the retrieval and generation heads are jointly optimized using the Gumbel-TopK trick (Kool et al., 2019), which mitigates the non-differentiability of the retrieval process.
Figure 4 illustrates the modified model architecture. The retrieval head generates a binary mask, $M\in\{0,1\}^{|C|}$, indicating which passages to select (1) or ignore (0). The selected passages are then passed to the generation head for response generation. The retrieval head consists of: (1) encoders for the query and passage, using the LCLM’s final hidden states, and (2) a scoring layer that computes relevance scores by concatenating their encoded vectors. The top $K$ passages are then selected for the generation head to produce a response.
Figure 4: A retrieval head is asdded to predict the Top-K context for generating final responses. The head is jointly optimized with the generation head.
More specifically, let $q$ denote a query and $C=\{c_{1},c_{2},\dots,c_{n}\}$ be the set of $n$ passages. For each pair $(q,c_{i})$, two single-layer encoders pool the final hidden states from an LCLM and generate a query vector, $\textbf{h}_{Q}$, and a passage vector, $\textbf{h}_{c_{i}}$:
$\textbf{h}_{q}=\text{enc}_{q}(q);\quad\textbf{h}_{c_{i}}=\text{enc}_{c}(c_{i}),$
Each pair $(\textbf{h}_{q},\textbf{h}_{c_{i}})$ is then concatenated to form an input vector $\textbf{v}_{i}=[\textbf{h}_{q};\textbf{h}_{c_{i}}]$. A single scoring layer then predicts a scalar relevance score $s_{i}$ for each pair, essentially computing $s_{i}=f(\textbf{v}_{i})$ where $f(\cdot)$ is our scoring function. With the set of scores $S=\{\dots s_{i}\dots\}$, we can finally identify the indices $T$ of the selected passages as the ones that receive the top $K$ scores,
$T=\text{TopK}(S,K).$
We then set the binary mask $M$ accordingly to filter the original context $C$ down to $C^{*}$, which is then used to prompt the LCLM to produce the final response.
5 Experiment
5.1 Setup
Benchmarks. We use LOFT (Lee et al., 2024) and our ICR2 benchmarks to evaluate LCLMs’ in-context retrieval and reasoning capabilities. LOFT tests retrieval, single- and multi-hop question answering, and reasoning using NaturalQuestions (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), and MuSiQue (Trivedi et al., 2022). ICR2 uses NaturalQuestions and HotpotQA, and additionally includes FEVER (Thorne et al., 2018) for fact verification and WoW (Dinan et al., 2018) for dialogue completion. Similar to LOFT, we report average scores across 100 test cases per task using the 32K context length versions, which is the maximum supported by all tested LCLMs.
Metrics. For the question answering and fact verification tasks, we use the exact match in (Lee et al., 2024; Adlakha et al., 2024). We use ROUGE (Lin, 2004) to assess on dialogue completion.
Training Details. We use ICR2’s training set to fine-tune all models. Specifically, we randomly sample 7500, 7500, 5000, and 5000 instances for NaturalQuestions, HotpotQA, FEVER, and WoW, respectively. To verify the effectiveness of our proposed methods, we focus our experiments on Mistral-Instruct-7B model (Jiang et al., 2023).
Baselines. We compare our approaches with the baselines in Sec. 3.3: Vanilla RAG, Closed-book, and Oracle RAG. We include the direct-answer SFT (SFT-DA) which follows Zhang et al. by concatenating the confounders with gold documents for SFT, and our methods in Sec. 4 include two retrieve-then-generate SFT variants — Retrieve-then-Answer (SFT-RTA), and Cite-Context-ID (SFT-CCI) — and the joint retrieval head training (RetHead) and retrieval attention probing (RAP).
5.2 Main Results
LOFT ICR2
NQ HPQA MUSI. NQ HPQA FEV WoW
GPT-4 w/
Close-book 0.58 0.57 0.28 0.61 0.50 0.85 0.16
Vanilla RAG 0.85 0.78 0.51 0.67 0.64 0.92 0.15
Oracle 0.88 0.87 0.72 0.79 0.82 0.94 0.18
Mistral-7B w/
Close-book 0.46 0.38 0.05 0.49 0.38 0.73 0.14
Vanilla RAG 0.64 0.62 0.27 0.53 0.50 0.81 0.13
Oracle RAG 0.89 0.81 0.41 0.83 0.81 0.94 0.18
[…]
SFT-Direct Answer 0.70 0.65 0.25 0.59 0.70 0.90 0.22
SFT-Retrieve-then-Answer 0.74 0.69 0.33 0.60 0.67 0.91 0.22
SFT-Cite-Context-ID 0.76 0.54 0.35 0.63 0.63 0.89 0.21
[…]
RetHead $w/\mathcal{L}_{gen}+\mathcal{L}_{ret}$ 0.15 0.13 0.07 0.48 0.54 0.9 0.21
RetHead $w/\mathcal{L}_{gen}$ 0.28 0.25 0.82 0.18
RetHead $w/\mathcal{L}_{ret}$ 0.39 0.44 0.82 0.13
[…]
SFT-DA w/ RAP 0.78 0.76 0.47 0.64 0.67 0.89 0.21
SFT-RTA w/ RAP 0.85 0.79 0.39 0.63 0.71 0.92 0.23
Table 4: Main results on LOFT (Lee et al., 2024) and ICR2 for our proposed methods applied on Mistral-2-7B. We also report the GPT-4 performance in the top panel. We highlight the improved and worsen performances compared with Vanilla RAG, and bold the best method.
As shown in Table 4, all SFT variants outperform the Vanilla RAG on both benchmarks, indicating that LCLMs struggle to effectively leverage context as a knowledge base for RAG tasks. Furthermore, the gap between SFT-DA and the Oracle RAG highlights that supervised fine-tuning alone is insufficient to achieve optimal results.
Among the SFT variants, SFT-RTA in general outperforms the others with an average improvement of 2%. Specifically on LOFT benchmark, both SFT-RTA and SFT-CCI outperform SFT-DA with 6% and 2% improvement on average, respectively, indicating the retrieve-then-generate strategy helps. On ICR2, however, all SFT variants perform the same. This demonstrates that ICR2 is a more discriminative benchmark than LOFT.
We apply RAP to all SFT models. On average, RAP enhances the models significantly, with SFT-DA + RAP improving by 6% and SFT-RTA + RAP by 8%. The best-performing approach, SFT-RTA + RAP, achieves notable gains on the challenging ICR2 benchmark, with improvements of 3%, 4%, 1%, and 1% on NaturalQuestions, HotpotQA, FEVER, and WoW, respectively. It also achieves top performance on 5 out of the 7 tasks, demonstrating its superiority.Remarkably, it achieves comparable performance with the state-of-the-art GPT-4-Turbo on LOFT and ICR2 while using a much smaller model. Finally, RAP decoding is more effective for SFT models than with the original model, as SFT better activates retrieval-specific attention heads for the approach (see Sec. 6.3).
For the joint retrieval head training (RetHead), we performed experiments on training only the generation head ($w/\mathcal{L}_{gen}$), only the retreieval head ($w/\mathcal{L}_{ret}$), and both ($w/\mathcal{L}_{gen}+\mathcal{L}_{ret}$). The last variant achieves the best performance, outperforming the Vanilla RAG baseline on ICR2. However, it is not comparable with the SFT variants.
5.3 In-Context Retrieval Performance
LCLM Methods LOFT ICR2 Avg.-All
NQ HPQA MUSI. Avg. NQ HPQA FEV. WoW Avg.
Mistral-2-7B SFT-RTA 0.81 0.83 0.37 0.67 0.58 0.78 0.64 0.43 0.61 0.63
SFT-CCI 0.65 0.61 0.26 0.51 0.63 0.73 0.69 0.49 0.64 0.58
[…]
RetHead $w/\mathcal{L}_{gen}$ 0.00 0.00 0.00 0.02 0.01 -
RetHead $w/\mathcal{L}_{ret}$ 0.33 0.49 0.51 0.49 0.46 -
[…]
SFT-DA + RAP 0.95 0.70 0.27 0.64 0.75 0.67 0.60 0.69 0.65 0.66
SFT-RTA + RAP 0.95 0.96 0.39 0.77 0.61 0.85 0.60 0.41 0.62 0.68
Table 5: Retrieval performance measured by recall rate for various methods using Mistral-2-7B.
Table 5.3 reports the recall rates for all methods, meaning how likely the models retrieve the relevant provenances. For the SFT variants, we evaluate the retrieval predictions produced in the retrieval phase. For the joint retrieval head training approach (RetHead), we analyze the predictions from the retrieval head. For RAP decoding, we assess the passages identified by the selected attention heads.
We find a strong correlation between a model’s recall rate and its downstream task performance, highlighting the importance of in-context retrieval ability for LCLMs. Our best approach, SFT-RTA + RAP, achieves the highest recall. In contrast, Vanilla RAG + RAP exhibits poor retrieval performance, consistent with its limited improvement on the downstream tasks. The near-random retrieval performance of joint training of retrieval head in LOFT explains its failure of generalization.
6 Discussion
6.1 Scaling the Supervised Fine-tuning
Figure 5: Scaling the number of training samples in ICR2’s training set.
We are also interested in how the performance of the proposed SFT variants scale with the training set size. We train the three variants, SFT-DA, SFT-RTA, and SFT-CCI, with the same 10K, 15K and 25K examples from ICR2’s training set, and report their performance on each task in the ICR2 benchmark, as shown in Fig 5. We observe that an increased training set size in general leads to an improved model performance. In particular, a smaller amount of training data fares worse with the retrieve-then-generate approaches, as they are by nature more challenging to learn compared to the SFT-DA approach, where answer is directly generated without an explicit retrieval step.
6.2 Blocking Context Attention in Retrieve-then-generate Model
ICR2 SFT-RTA Blocking Context
Metrics Recall Metrics Recall
NQ 0.6 0.58 0.61 (.$\uparrow.01$) 0.59 (.$\uparrow.01$)
HPQA 0.67 0.78 0.68 (.$\uparrow.01$) 0.78 (=.$=.00$)
FEVER 0.91 0.64 0.91 (=.$=.00$) 0.66 (.$\uparrow.02$)
WoW 0.22 0.43 0.21 (.$\downarrow.01$) 0.46 (.$\uparrow.03$)
[…]
Table 6: SFT-RTA’s performance before and after blocking the contextual knowledge base with attention mask.
To verify if models actually learn to generate the final responses only from the retrieval predictions in our retrieve-then-generate methods (Sec. 4.1), additional experiments were performed where we block attention to be paid onto the context beyond the retrieval predictions at the generation step, i.e.,
$p(y\mid q,c)=\sum_{z_{i}\in Z}p(y\mid q,z_{i})p(z_{i}\mid q,c),$ (6)
Table 6.2 reports the impact of the blocking on the SFT-RTA variant. We observe the performance essentially stays the same, indicating the model indeed learned to generate largely on the retrieved passages, ignoring the original context.
6.3 Effect on Attention Heads of Retrieve-then-Generate Fine-tuning
Figure 6: Attention heads with above-zero hit rates. SFT produces more retrieval attention heads. Retrieve-then-generate training activates a higher peak of hit rate.
To understand if fine-tuning sharpens attention heads’ focus on relevant passages, we compare the hit rates (Sec. 4.2) achieved by the attention heads between Vanilla RAG and all of our SFT variants (Sec. 4.1), and the results are shown in Figure 6.
Similar to (Wu et al., 2024), we find that a small group of attention heads can obtain higher hit rates than the others. However, unlike Vanilla RAG, SFT methods produce more retrieval-focused attention heads, and achieve higher peak hit rates. This demonstrates the effectiveness of our SFT methods and our curated ICR2 training set in enhancing LCLMs’ performance of in-context retrieval.
Among the three SFT variants, we find that SFT-RTA in general achieves higher peak hit rates and activates more attention heads for retrieval. In particular, SFT-CCI is not as effective in recruiting as many attention heads, possibly because context IDs themselves are not informative enough for the model to learn the retrieval task well. We also note that SFT-DA fares a lot worse on FEVER, possibly due to the lack of chain-of-thought style of assist.
6.4 Retrieval Attention Probing
Figure 7: Effect of adjusting $Q$ (number of attention heads for retrieval) and $M$ (number of passages to select) when applying SFT-RTA + RAP. Left column: average model performance measured by the tasks-specific metrics. Middle column: attention heads’ retrieval measured by the recall rate. Right column: RTA’s retrieval measured by the exact match.
Our inference-time method RAP (Sec. 4.2) uses two hyperparameters: $Q$ is the number of the attention heads we recruit for retrieval, and $M$ is the number of passages each head retrieves. In this section, we apply different value settings when deploying the SFT-RTA + RAP combined approach to explore their effect on model performance.
Figure 7 shows the high-level results on both LOFT and ICR2 (Appendix B has more details). As expected, we observe in the middle column that an increasing $M$ or $Q$ will increase the recall rate as the resulting larger pools of selected passages will more likely include the relevant ones. We also find that increasing $M$ as opposed to $Q$ is more effective in improving the recall rate as certain task such as HotpotQA requires multi-hop retrieval. However, the increased recall rate does not always translate to a better performance (left column) or RTA’s recall rate , as a higher $M$ or $Q$ may also introduce more confounders. This suggests the further reduction in confounding effects is still an opening future work.
6.5 Decoding Speed
Benchmark SFT-DA SFT-DA + RAP
NQ 4.23 4.99 (.$\uparrow.76$)
HotpotQA 4.19 4.81 (.$\uparrow.62$)
MUSIQUE 4.24 5.05 (.$\uparrow.81$)
Table 7: Latency for SFT-DA with and without the RAP decoding. Time is measured in seconds per query.
Our final analysis is on the efficiency of RAP. Based on the SFT-DA variant, we report its latency with and without the RAP enhancement in Table 7. We find that the RAP decoding does not increase latency significantly, despite it adds one additional decoding step to the base method. This can be attributed to the much shorter context retrieved by the attention heads, thus avoiding expensive long-context computation as in the baseline. This increase can be further reduced with approaches such as KV caching, which we leave for future work.
7 Conclusion
In this paper, we introduce ICR2, a new benchmark designed as a more realistic and discriminative benchmark for evaluating LCLM’s abilities in in-context retrieval and reasoning. Our findings highlight the limitations for the current models. Additionally, we propose three methods —retrieve-then-generate fine-tuning, retrieval attention probing, and joint retrieval head training — to enhance models, achieving the results comparable to GPT-4, but with a smaller model footprint.
Ethical Considerations
We do not expect any ethical concerns to be raised with respect to this work.
Limitations
We acknowledge several limitations in this work. First, most experiments were conducted with a context length of 32K tokens. Our findings indicate that while many LCLMs claim to support longer contexts, their performance on tasks with 32K tokens remains suboptimal. Future work could focus on extending ICR2 and the proposed approaches to effectively support scenarios with longer context lengths.
Second, while Joint Retrieval Head Training demonstrates improved performance compared to Vanilla RAG, it still falls short of the performance achieved by the SFT variant. Future research could explore improved architectural designs to better integrate the supervision signals from both the retrieval and generation tasks.
Finally, our evaluation primarily utilizes the Mistral-7B model due to computational constraints. Extending the proposed methods to other LCLMs would provide a broader assessment of their generalization capabilities and effectiveness across different model architectures.
Appendix A Prompt Template for the Retrieval-augmented Generation
Task Prompt Template
QA [INST] Please answer the following question given the following passages:
{Corpus}
Question: {Query}
Answer: [/INST]
Fact Verification [INST] According to the following passages, please verify the given claim
and predict your judgment on its factuality as TRUE or FALSE:
{Corpus}
Claim: {Query}
Judgement: [/INST]
Dialogue Completion [INST] According to the given passages, please provide a single response to
complete the following conversation by role-playing as either Person A or
Person B. Your response should be as knowledgeable and coherent with the
conversation history as possible:
{Corpus}
Conversation: {Query}
[/INST]
Table 8: Prompt template used for RAG tasks in our experiments. {Corpus} refers to the provided contextual knowledge base, and {Query} refers to a query in ICR2 or LOFT. A {Query} can be a question in the question answering tasks (Kwiatkowski et al., 2019; Yang et al., 2018), a claim to be verified in the fact verification task (Thorne et al., 2018), or a conversation history in the dialogue completion task (Dinan et al., 2018).
We show the Corpus-in-Context (CiC; Lee et al. 2024) prompt template used in our experiments in Table 8.
Appendix B Detailed Results for RAP Decoding
We show the detailed results for each task of LOFT and ICR2 in Figure 8 and 9.
Figure 8: Effect of adjusting $Q$ (number of attention heads for retrieval) and $M$ (number of passages to select) when applying SFT-RTA + RAP on all tasks in LOFT. Left column: average model performance measured by the tasks-specific metrics. Middle column: retrieval performance measured by the recall rate. Right column: retrieval performance measured by the exact match.
Figure 9: Effect of adjusting $Q$ (number of attention heads for retrieval) and $M$ (number of passages to select) when applying SFT-RTA + RAP on all tasks in ICR2. Left column: average model performance measured by the tasks-specific metrics. Middle column: retrieval performance measured by the recall rate. Right column: retrieval performance measured by the exact match.
Appendix C Details of Experiment Setup
C.1 Training Details
For all models, we set the base value of the rotary position embedding to $61e$ 6 following Su et al. (2024). Training is conducted with a batch size of 1 per Nvidia A100-40G GPU, and gradients are accumulated every 4 steps. Each model is trained for up to 10,000 steps (approximately 2 epochs), with the best-performing checkpoint on the validation set selected as part of an early-stopping strategy. The learning rate is set to $51e-$ 5, and the maximum sequence length during training is capped at 32,768 tokens, discarding any sequences exceeding this threshold.
C.2 Inference Parameters
During inference, we use greedy decoding for all models, allowing a maximum generated sequence length of 1,024 tokens. For RAP decoding, 100 random instances from the validation set are used to probe the attention heads responsible for retrieval. For SFT-DA, retrieval is performed using the designated retrieval attention heads based on the first and only decoded token. For SFT-RTA, retrieval is performed using all tokens generated in the retrieval step to ensure complete context coverage.
C.2.1 RAP Hyperparameter Settings
In this section, we detail the hyperparameters used in applying RAP decoding to SFT-DA and SFT-RTA on ICR2 and LOFT.
For SFT-DA on ICR2:
  • NaturalQuestions: $Q=4,M=1$
  • HotpotQA: $Q=8,M=1$
  • FEVER: $Q=4,M=4$
  • WoW: $Q=4,M=4$
For SFT-DA on LOFT:
  • NaturalQuestions: $Q=4,M=4$
  • HotpotQA: $Q=4,M=4$
  • WoW: $Q=8,M=2$
For SFT-RTA on ICR2:
  • NaturalQuestions: $Q=2,M=1$
  • HotpotQA: $Q=2,M=1$
  • FEVER: $Q=2,M=8$
  • WoW: $Q=4,M=8$
For SFT-RTA on LOFT:
  • NaturalQuestions: $Q=8,M=4$
  • HotpotQA: $Q=4,M=2$
  • WoW: $Q=8,M=2$
Appendix D Effect of Retrieval Delineation
Benchmark SFT-RTA $w/o$ <RET> SFT-CCI $w/o$ <RET>
LOFT NQ 0.74 0.7 (.$\downarrow.04$) 0.76 0.63 (.$\downarrow.13$)
HPQA 0.69 0.66 (.$\downarrow.03$) 0.54 0.58 (.$\uparrow.04$)
MUSI. 0.33 0.24 (.$\downarrow.09$) 0.35 0.24 (.$\downarrow.11$)
ICR2 NQ 0.60 0.56 (.$\downarrow.04$) 0.63 0.61 (.$\downarrow.02$)
HPQA 0.67 0.7 (.$\downarrow.03$) 0.63 0.65 (.$\uparrow.02$)
FEVER 0.91 0.92 (.$\uparrow.01$) 0.89 0.86 (.$\downarrow.03$)
WoW 0.22 0.20 (.$\downarrow.02$) 0.21 0.22 (.$\uparrow.01$)
Table 9: Effect of removing the special boundary tokens for the retrieve-then-generate variants.
As outlined in Section 4.1, our retrieve-then-generate variants, SFT-RTA and SFT-CCI, utilize special symbols <RETRIEVAL> and </RETRIEVAL> to explicitly mark the boundaries of the retrieval step. To examine the impact of these boundary tokens, we conducted an ablation study. As shown in Table 9, removing these tokens in general leads to a decline in model performance. This suggests that clearly delineating the retrieval step is crucial for the model to learn effectively.